OpenStudy (anonymous):

7. will give medal Find the value for x that makes the equation -3(x + 10) + 2x = 75 true. plzz show me how to do this

OpenStudy (anonymous):

@AskMeLotss

OpenStudy (anonymous):

im sorry but I gtg... :(

OpenStudy (anonymous):

bye @AskMeLotss

OpenStudy (anonymous):

(-3x-30)+2x=75 -3x+2x-30=75 -x-30=75 -x=75+30 -x=105 x=-105

OpenStudy (anonymous):

thanks @ArifL

OpenStudy (anonymous):

no prob